# GraphTeam: Facilitating Large Language Model-based Graph Analysis via Multi-Agent Collaboration

### v2

@@ 4.3. Ablation Study (RQ2)
To evaluate the contribution of each component, we conduct a systematic ablation study. We assess the impact of Input-Output Normalization by removing the question or answer agent. We examine the influence of Knowledge Retrieval components in the search agent, specifically documentation and experience. Additionally, the effectiveness of Problem Solving designs are evaluated by eliminating the coding or reasoning agent, as well as the retry mechanism from the coding agent. This comprehensive ablation enables us to quantify each component’s significance to the system performance. The overall results are presented in Tab.1 and 2.
@@
Input-Output Normalization: For most benchmark problems, the question agent significantly contributes (25.44%) by breaking the problem into individual components. For a few benchmark such as GraphInstruct (Luo et al., 2024), where the structure of the question is already clear and straightforward, the impact of the question agent is less noticeable (less than 2.28%). The answer agent’s correction of the format is also an essential part of correctly answering the questions (0.54%- 3.33%).
@@
External Knowledge Retrieval: The search agent plays a significant role. When solving graph analysis problems, providing the large language model with prior knowledge of related problem-solving experiences or API documentation can help the model solve the problem more effectively. The impact of experience is greater than that of documentation, indicating that LLMs solve problems in a way similar to humans with an average improvement of 24.63%. They can derive better solutions from prior experiences, effectively addressing similar problems. Documentation can enhance the model’s programming capabilities. When relevant experiences are not available, documentation is still needed to learn graph analysis methods and subsequently solve corresponding problems (0.64% to 7.1%).
@@
Problem Solving: The coding agent plays a crucial role. Without programming and relying solely on reasoning, the system’s performance would significantly decrease, indicating that programming is an effective approach for solving graph analysis problems (27.05% to 58.50%). The retry mechanism does not always have a significant overall effect, but it shows clear benefits for hard problems with a maximum improvement of 57.87%. This indicates that Hard problems require more complex code, which cannot be written correctly in a single attempt and needs multiple tries to solve effectively. The reasoning agent does not have a significant impact on problems where programming achieves good results, but it plays an important role in problems where programming is insufficient and reasoning is required (0.81% to 34.44%).
@@
The contribution of each component of the system varies across different benchmarks. Since the core problem-solving component of this system is the coding agent, programming to solve problems is a crucial and effective approach for graph analysis. For GNNs, in particular, coding is an indispensable component. Therefore, removing the coding agent leads to a significant drop in performance for all benchmarks (42.24% in average). Some benchmarks do not have strict requirements for the answer format (NLGraph, GraphWiz, and LLM4DyG), so removing the answer agent does not significantly impact the overall system performance (0.76% in average). Since Talk like a Graph contains simple questions, most problems can be solved correctly through coding without the need for the reasoning agent, hence the benchmark has minimal need for the reasoning agent (6.36%). Since GNN-AutoGL lacks documentation information, this module does not affect system performance (1.56%). Similarly, GraphInstruct’s problem-solving does not rely on documentation, so documentation has little impact on the system’s performance on this benchmark (0.04%).
Figure 3. Performance with respect to different task categories.
Figure 4. Performance with respect to different output formats.
Figure 5. Hyper-parameter analysis of four hyper-parameter in the proposed GraphTeam.
4.4. Analysis on Different Groupings (RQ3)
To provide a more nuanced understanding of GraphTeam’s capabilities and limitations, we conduct a detailed analysis of its performance across various task categories and output formats. This fine-grained approach allows us to identify specific strengths and weaknesses.
Group by Task Categories. We propose a categorization scheme that covers a broad spectrum of graph-related problems. We classify the questions into four categories based on the nature of the task: (1) basic: fundamental graph understanding, (2) algorithm-macro-level: analysis of entire graph structures, (3) algorithm-micro-level: local properties of nodes and edges, and (4) GNN: graph neural networks. As shown in Fig. 3, GraphTeam achieves an accuracy of over 90% on both basic and macro-level algorithms, and 85.2% on micro-level tasks, indicating that the system can effectively handle common graph analysis tasks. However, all systems, including GraphTeam, struggle with GNN-related tasks, suggesting a need for improvement in this area.
These results reveal interesting insights: For micro-level tasks, removing the reasoning agent causes a significant 21.2% drop in performance, highlighting the complexity of these problems and the limitations of the coding agent alone. In GNN-related tasks, while the coding agent and experience are important, the question agent plays a crucial role (38.8% impact when removed). This underscores the importance of effectively extracting key information from GNN problem statements. These findings point to areas for focused improvement, particularly in enhancing reasoning capabilities for micro-level tasks and information extraction for GNN-related problems.
Group by Output Formats. To account for the diversity in problem outputs, we categorize the questions based on their required response formats: (1) yes/no, (2) digits, (3) list/set, and (4) others. This classification allows us to assess system performance across different types of graph analysis outputs. As shown in Fig. 4, GraphTeam demonstrates strong performance on yes/no, digits, and list/set outputs, achieving an accuracy of over 85% for each of these categories. However, the system’s performance drops notably for more complex output formats, with an accuracy of only 70.1%. This decline in accuracy for the “others” format indicates that GraphTeam, particularly its answer agent, needs enhancement in handling advanced data structures and more complex output formats. Improving the system’s capability to generate and structure complex outputs could significantly boost its overall effectiveness in addressing a wider range of graph analysis problems.
4.5. Hyper-parameter Sensitivity (RQ4)
@@
We conduct a study on key hyper-parameters using the NLGraph (Wang et al., 2024a) benchmark to determine optimal settings or optimization strategies. Fig. 5 presents our analysis of several important hyper-parameters. For the number of candidate experiences $N_{exp}$ in knowledge base construction, our experiments show that larger $N_{exp}$ values correlate with better system performance, as they provide more candidate experiences for problem-solving assistance. The similarity matching threshold $\delta$ between experiences and documents achieves optimal problem-solving performance at $0.85$, enabling correct matching of similar experiences when available and document retrieval when experiences are absent. Increasing the maximum retries of the coding agent and the number of self-checking iterations of the answer agent improves system performance, but with diminishing returns as these values increase, eventually plateauing. These findings provide valuable insights for optimizing the system’s configuration and balancing performance gains against computational costs.
5. Conclusion
@@
To address the limitations of previous LLM-based graph analysis methods, we draw inspiration from human problem-solving processes, and design a multi-agent system named GraphTeam based on LLMs. GraphTeam consists of five agents from three key modules, namely input-output normalization, external knowledge retrieval, and problem solving. The agents can cooperate with each other for solving complex graph analysis problems. Extensive experiments on six benchmarks demonstrate that GraphTeam achieves state-of-the-art results with an average improvement of 25.85% in terms of accuracy. For future work, we will consider to fine-tune the LLMs in GraphTeam for better performance, and further optimize the system pipeline to accommodate more tasks.
Appendix A Prompt Templates
A.1. Question Agent
Requirement Analyst
Here is the task: task.
@@ A.1. Question Agent
As the assistant_role, you need to extract key information from the task, specifically focusing on the following elements:
@@
Input(str): This refers to the graph data or graph file required by the task. You must extract **all relevant information** about the graph data, including but not limited to graph representation (adjacency list, adjacency matrix, edge list, etc.), node and edge details, and any other relevant information.
Problem(str): This is the specific problem that needs to be solved.
@@
A.2. Search Agent
Research Assistant
You are tasked as a assistant_role to retrieve relevant information for the task: query.
@@ A.2. Search Agent
A.3. Coding Agent
Graph Learning Specialist
Based on the available search result: search_result, your goal is to write Python code that solves the specified graph learning problem using libraries like NetworkX.
@@ A.3. Coding Agent
Input: input,
Problem: problem,
Output Format: output_format.
@@
A.4. Reasoning Agent
Graph Learning Expert
As a assistant_role, your task is to analyze and solve a graph analysis problem.
@@ A.4. Reasoning Agent
Here is the task based on the extracted information: Input: input, Problem: problem, Output Format: output_format.
@@
’###output:
solution
A.5. Answer Agent
Output Format Specialist
As a assistant_role, your role is to ensure that the generated output strictly adheres to the specified format requirements.
You will be provided with both the required output format (example format) and the actual output. Your task is to compare the two and determine if the actual output matches the format requirements. If the output does not meet the format, you must adjust the format accordingly while **preserving every part of the original content, meaning, and intent without deleting or changing any sections**.
Important: The required format provided is an **example** that illustrates the expected structure, not the actual content or final answer. Do **not** alter the original meaning, sections, sentences, or intent of the actual output. Your task is only to adjust the format if necessary.
When adjusting the format, ensure that **no parts are removed, shortened, or reworded in a way that changes the original meaning or omits content**. The adjustment should focus on structural changes such as spacing, indentation, or organization.
Once the adjustment is made, compare the adjusted output with the original to validate that **all sections, sentences, and words are fully preserved** in the final output.
@@ A.5. Answer Agent
- **Required format (example)**: output_format
- **Actual output** (content to be reviewed): output
Carefully review both and ensure that the actual output conforms to the structure of the example format, while **maintaining the original content, meaning, sections, and intent**.
Note: Only adjust the **format** if necessary. Do **not** change the actual content, meaning, or intent of the output. If the format is already correct, no adjustment is needed.
@@
- **need_adjustment** (True/False): Indicating whether an adjustment is required.
- **output** (str): If no adjustment is needed, this should match the original output exactly. If an adjustment is needed, this should contain the adjusted output, with only format changes applied, without altering the content, meaning, sections, or intent.
Appendix B GNN-AutoGL Benchmark Construction
For the GNN-AutoGL benchmark, we use a unified template for construction. Specifically, we ask the large language model (LLM) to write code that constructs a GNN model using AutoGL based on the given prompt. Given the numerous model parameters, specifying each parameter and requiring the LLM to write corresponding model code is challenging. Therefore, we select only some parameters as requirements and evaluate whether the LLM correctly includes them. For a problem with $K$ specified parameters, if the code runs successfully, we use a Python script to assess the correctness of the $k$ specified parameters, resulting in a score of $k/K$. If the code fails to run, the score is zero.
Below is a template for constructing a problem.
In graph theory, several fundamental tasks are crucial for various applications.
Shortest Path algorithms, like Dijkstra’s and Bellman-Ford, help find the minimum distance between nodes in a weighted graph, playing a key role in routing and network design.
Hamiltonian Cycle detection determines whether a cycle exists that visits each vertex exactly once, a problem common in scheduling and logistics.
Subgraph Matching focuses on finding a smaller graph pattern within a larger graph, widely used in chemical informatics and pattern recognition.
These tasks, along with Graph Coloring, Spanning Trees, and Maximum Flow, demonstrate the diversity and utility of analyzing graph structures.
With the rise of Graph Neural Networks (GNNs), new ways of addressing graph-based tasks, like node classification, link prediction, and clustering, have emerged, allowing models to learn from the inherent relationships in graph-structured data.
Task: Given the dataset name dataset, can you use Python library to train a layer number-layer GNN model name model with hidden dimension as hidden dimension, dropout as dropout and predict the accuracy ?
A Python code example for solving GNN-AutoGL problems.
import pkg_resources
@@ Appendix B GNN-AutoGL Benchmark Construction
import torch
from autogl.module.train import Acc
from autogl.solver.utils import get_graph_labels,
get_graph_masks
from autogl.datasets import build_dataset_from_name
cora_dataset = build_dataset_from_name(’cora’)
@@
    ’num_layers’: 2,
    ’hidden’: [128, 256],
    ’dropout’: 0.6,
    ’act’: ’relu’,
    ’eps’: ’false’

### v3

@@ 4.3 Ablation Study
To evaluate the contribution of each component, we conduct a systematic ablation study of GraphTeam using GPT-4o-mini. We assess the impact of Input-Output Normalization by removing the question or answer agent. We examine the influence of Knowledge Retrieval components in the search agent, specifically documentation and experience. Additionally, the effectiveness of Problem Solving designs are evaluated by eliminating the coding or reasoning agent, as well as the retry mechanism from the coding agent. This comprehensive ablation enables us to quantify each component’s significance to the system performance. The results are presented in Table 2.
Table 2: Performance comparison between GraphTeam and ablation variants in terms of accuracy (%).
Category Methods NLGraph Talk like a Graph GraphInstruct GraphWiz LLM4DyG GNN-AutoGL
GraphTeam (GPT-4o-mini) 97.8 94.5 88.1 88.6 94.8 48.4
     - question agent 79.2 82.1 85.9 62.8 58.6 14.8
Input-Output Normalization      - answer agent 96.7 93.9 85.3 88.0 94.2 45.8
     - documentation 96.4 93.8 86.1 88.0 94.0 46.9
Knowledge Retrieval      - experience 54.7 92.0 76.6 82.8 60.3 4.8
     - coding agent 39.3 60.7 61.1 46.4 59.3 -
     - retry 87.7 93.5 85.7 83.5 88.5 44.3
Problem Solving      - reasoning agent 96.9 93.9 85.5 54.2 93.7 46.4
@@
Input-Output Normalization: For most benchmark problems, the question agent significantly contributes (25.3%) by breaking the problem into individual components on average. For a few benchmark such as GraphInstruct Luo et al. (2024), where the structure of the question is already clear and straightforward, the impact of the question agent is less noticeable (less than 2.2%). The answer agent’s correction of the format is also an essential part of correctly answering the questions (0.6-2.8%).
@@
External Knowledge Retrieval: The search agent plays a significant role. When solving graph reasoning problems, providing LLM with related problem-solving experiences or API documentation can help the model solve the problem more effectively. The impact of experience is greater than that of documentation, indicating that LLMs solve problems in a way similar to humans with an average improvement of 23.5%. They can derive better solutions from prior experiences, effectively addressing similar problems. When relevant experiences are not available, documentation can also help enhance the model’s programming capabilities in solving corresponding problems (0.6% to 2.0%).
@@
Problem Solving: The coding agent plays a crucial role. Without programming and relying solely on reasoning, the system’s performance would significantly decrease, indicating that programming is an effective approach for solving graph reasoning problems (27.0% to 58.5%). The retry mechanism also brings a positive effect (1.0% to 10.1%) by debugging programs. The reasoning agent does not have a significant impact on problems where programming achieves good results, but it plays an important role in problems where programming is insufficient and reasoning is required (0.9% to 34.4%).
@@
The contribution of each component of the system varies across different benchmarks. Since the core problem-solving component of this system is the coding agent, programming to solve problems is a crucial and effective approach for graph reasoning. For GNNs, in particular, coding is an indispensable component. Therefore, removing the coding agent leads to a significant drop in performance for all benchmarks (40.9% on average). Some benchmarks do not have strict requirements for the answer format (Talk like a graph, GraphWiz, and LLM4DyG), thus removing the answer agent does not significantly impact the overall system performance. Since questions in Talk like a Graph are relatively simple, most problems can be correctly solved through coding without the need for the reasoning agent. Hence this benchmark has minimal need for the reasoning agent. In summary, the full model always perform the best, validating the necessity of each design.
4.4 Analysis on Different Groupings
Group by Task Categories. We classify the questions into four categories based on the nature of the task: (1) basic: fundamental graph understanding, (2) macro-level: reasoning over entire graph structures, (3) micro-level: local properties of nodes and edges, and (4) GNN: graph neural networks. As shown in Figure 3, GraphTeam (GPT-4o-mini) achieves over 90% accuracies on the first two categories, and 85.6% on micro-level tasks, ranking the first over all four groups. While GraphTeam (Qwen2.5-Coder-7B) always ranks the second, achieving over 76% on the first three categories. These results indicate that our framework can effectively handle common graph reasoning tasks. For most categories (basic, micro-level, and GNN), we observe consistent improvements when utilizing agent frameworks. But on macro-level tasks, we observe that AutoGPT, Open Interpreter, and AutoGen perform worse than using GPT-4o-mini alone. We find that this degradation is caused by misunderstanding question requirements during long reasoning processes. Thus, employing the question agent to explicitly identify key information is necessary for graph reasoning. Besides, all systems struggle with GNN-related tasks, suggesting a need for improvement in future work.
Figure 3: Performance with respect to different task categories.
Group by Output Types. We categorize problems according to their required data formats: true/false questions (yes/no), numerical calculations (digits), analytical questions (list/set), and other types, with results presented in Figure 4. Similarly, GraphTeam with GPT-4o-mini and Qwen2.5-Coder-7B always rank the first and second across all four output formats. In particular, GraphTeam (GPT-4o-mini) reaches or approaches 90.0% for the first three categories, showing a strong generalization ability. The base LLMs perform good for true/false questions, but are inferior to agent-based methods for problems requiring more complex output formats.
Figure 4: Performance with respect to different output formats.
Figure 5: Average accuracies (%) and token consumptions of different agent-based methods. While maintaining comparable token consumption with typical agent frameworks, GraphTeam substantially improves the average accuracy for graph reasoning.
4.5 Efficiency Analysis
To verify whether our proposed GraphTeam is cost effective, we present average accuracy and token consumption for all agent-based methods based on GPT-4o-mini in Figure 5. GraphTeam achieves an average accuracy of 85.4% with relatively low token consumption. The token cost of GraphTeam is competitive with single-agent baselines, and is significantly lower than ChatDev. Though ChatDev demonstrates notable accuracy advantages over other agent frameworks, it consumes approximately 5.6 times more tokens than GraphTeam. This indicates that our framework design is better optimized for graph reasoning tasks, achieving promising performance with acceptable resource consumption.
4.6 Hyper-parameter Analysis
@@
As shown in Figure 6, we conduct experiments to study the influence of key hyper-parameters using the NLGraph Wang et al. (2024a) benchmark. For the number of candidate experiences $N_{exp}$ in knowledge base construction, our experiments show that larger $N_{exp}$ values correlate with better system performance, as they provide more candidate experiences for problem-solving assistance. The similarity matching threshold $\delta$ between experiences and documents achieves optimal problem-solving performance at $0.85$, enabling correct matching of similar experiences when available and document retrieval when experiences are absent. Increasing the maximum retries of the coding agent and the number of self-checking iterations of the answer agent improves system performance, but with diminishing returns as these values increase, eventually plateauing. These findings provide valuable insights for optimizing the system’s configuration and balancing performance gains against computational costs.
Figure 6: Analysis of key hyper-parameters in the proposed GraphTeam.
5 Conclusion
@@
In this work, we draw inspiration from human problem-solving processes, and design a multi-agent system named GraphTeam for graph reasoning. GraphTeam consists of five agents across three key modules, namely input-output normalization, external knowledge retrieval, and problem solving. The agents can cooperate with each other for solving complex graph reasoning problems. Extensive experiments show that GraphTeam achieves state-of-the-art results with impressive accuracy improvements.
As the first work to address LLM-based graph reasoning problems using a multi-agent system, our workflow prioritizes simplicity, with most agents operating sequentially. For future work, we will consider to explore more sophisticated communications between agents, and automatically optimize the system pipeline to accommodate more tasks. It is also possible to fine-tune the LLMs in GraphTeam for better performance.
Appendix A Method Details
A.1 Problem-Solving Experience Collection
[…]
Algorithm 2 Problem-Solving Experience Collection
0:  Training dataset $\mathcal{D}_{\text{train}}$, validation dataset $\mathcal{D}_{\text{valid}}$, candidate experience size $N_{exp}$ for each problem type $t\in\mathcal{T}$.
0:  Experience knowledge base $\mathcal{K}_{exp}$.
1:  for each problem $(q,a,t)$ from $\mathcal{D}_{\text{train}}$ do
2:     if $|\mathcal{K}_{exp}^{t}|$ $<$ $N_{exp}$ then
[…]
4:        if answer equals to ground truth $a$ then
[…]
6:        end if
7:     end if
8:  end for
9:  for each problem $(q,a,t)$ from $\mathcal{D}_{\text{valid}}$ do
10:     for $e\in\mathcal{K}_{exp}^{t}$ do
[…]
12:        if answer equals to ground truth $a$ then
13:           $\text{Utility}_{e}\leftarrow\text{Utility}_{e}+1$
14:        end if
15:     end for
16:  end for
17:  for $t\in\mathcal{T}$ do
18:     $\mathcal{K}_{exp}^{t}\leftarrow\text{argmax}_{e\in\mathcal{K}_{exp}^{t}}\text{Utility}_{e}$
19:  end for
20:  $\mathcal{K}_{exp}\leftarrow\{\mathcal{K}_{exp}^{t}\}_{t\in\mathcal{T}}$
A.2 Prompt Templates
Below are the prompts of the five agents used in GraphTeam.
A.2.1 Question Agent
Here is the task: {original question}.
@@ A.2.1 Question Agent
As a requirement analyst, you need to extract key information from the task, specifically focusing on the following elements:
Reformatted_Problem (str): This is the specific problem that needs to be solved.
@@
Input_Data (str): This refers to the graph data or graph file required by the task. You must extract all relevant information about the graph data, including but not limited to graph representation (adjacency list, adjacency matrix, edge list, etc.), node and edge details, and any other relevant information.
@@
A.2.2 Search Agent
You are tasked as a research assistant to retrieve relevant information for the task: {reformatted_question, graph_type}.
@@ A.2.2 Search Agent
A.2.3 Coding Agent
Based on the available search result: {konwledge of the question}, your goal is to write Python codes that solve the specified graph learning problem using libraries like NetworkX.
@@ A.2.3 Coding Agent
Input_Data: input_data,
Output_Format: output_format.
@@
A.2.4 Reasoning Agent
As a graph learning expert, your task is to analyze and solve a graph reasoning problem: {refomatted_question}.
@@ A.2.4 Reasoning Agent
Here is the task based on the extracted information: Input_Data: input_data, Reformatted_Problem: reformatted_problem, Output_Format: output_format.
@@
A.2.5 Answer Agent
As an assistant_role, your role is to ensure that the generated output strictly adheres to the specified format requirements.
You will be provided with both the required output format (example format) and the actual output. Your task is to compare the two and determine if the actual output matches the format requirements. If the output does not meet the format, you must adjust the format accordingly while preserving every part of the original content, meaning, and intent without deleting or changing any sections.
Important: The required format provided is an example that illustrates the expected structure, not the actual content or final answer. Do not alter the original meaning, sections, sentences, or intent of the actual output. Your task is only to adjust the format if necessary.
When adjusting the format, ensure that no parts are removed, shortened, or reworded in a way that changes the original meaning or omits content. The adjustment should focus on structural changes such as spacing, indentation, or organization.
Once the adjustment is made, compare the adjusted output with the original to validate that all sections, sentences, and words are fully preserved in the final output.
@@ A.2.5 Answer Agent
- Required format (example): output_format
- Actual output (content to be reviewed): output
Carefully review both and ensure that the actual output conforms to the structure of the example format, while maintaining the original content, meaning, sections, and intent.
Note: Only adjust the format if necessary. Do not change the actual content, meaning, or intent of the output. If the format is already correct, no adjustment is needed.
@@
- need_adjustment (True/False): Indicating whether an adjustment is required.
- output (str): If no adjustment is needed, this should match the original output exactly. If an adjustment is needed, this should contain the adjusted output, with only format changes applied, without altering the content, meaning, sections, or intent.
Appendix B Details about Datasets and Metrics
B.1 Dataset Statistics
We list the statistics of the six benchmarks in Table 3, including task categories, output formats, maximum graph sizes, and the number of samples in experience and test sets.
Table 3: Statistics of six graph reasoning benchmarks.
Task Category Output Format Graph Size Total Experiences Test Set Size
NLGraph Macro-level, Micro-level, GNN yes/no, digits, list/set, others $\sim 10^{1}$ 80 1,000
Talk like a Graph Basic, Macro-level yes/no, digits, list/set $\sim 10^{1}$ 80 8,000
GraphInstruct Basic, Macro-level, Micro-level yes/no, digits, list/set $\sim 10^{1}$ 170 5,100
GraphWiz Macro-level, Micro-level yes/no, digits, list/set $\sim 10^{2}$ 90 3,600
LLM4DyG Basic, Micro-level yes/no, digits, list/set $\sim 10^{1}$ 90 2,700
GNN-AutoGL GNN others $\sim 10^{3}$ 10 200
B.2 GNN-AutoGL Benchmark Construction
For the GNN-AutoGL benchmark, we use a unified template for construction. Specifically, we ask the large language model (LLM) to write code that constructs a GNN model using AutoGL based on the given prompt. Given the numerous model parameters, specifying each parameter and requiring the LLM to write corresponding model code is challenging. Therefore, we select only some parameters as requirements and evaluate whether the LLM correctly includes them. For a problem with $K$ specified parameters, if the code runs successfully, we use a Python script to assess the correctness of the $k$ specified parameters, resulting in a score of $k/K$. If the code fails to run, the score is zero. Below is an example of problems in this benchmark.
Given the cora dataset, can you use autogl to train a $2$-layer $gcn$ model with hidden dimension as $[128,256]$, dropout as $0.6$ and predict the accuracy ?
A Python code example for solving the above problem is shown below.
@@ B.2 GNN-AutoGL Benchmark Construction
from autogl.datasets import
build_dataset_from_name
cora_dataset =
build_dataset_from_name(‘cora’)
@@
    ‘num_layers’: 2,
    ‘hidden’: [128, 256],
    ‘dropout’: 0.6,
    ‘act’: ’relu’,
    ‘eps’: ’false’